1) Let  be a function from Z to R, such that , then  is

a) an increasing function   

b) a strictly increasing function

c) a decreasing function

d) an onto function


1
Expert's answer
2021-05-07T09:55:26-0400

Let f be a function from Z to R, such that f(x)=x/10

So, "f(x)=\dfrac{x}{10}"


Graph of f(x) in its domain:




Now, Interval in which f(x) to be determined is ("-\infty,\infty" )

So, a= "-\infty" and b= "\infty"

and f(a)= "-\infty" and f(b)= "\infty"

Hence f(a)< f(b)

Now, f'(x)= "\dfrac{1}{10}>0\ \ \ \forall \ \ x\in(-\infty,\infty)"


So, it is clear that  f(a)<f(b) and f'(x)>0

Hence f(x) is Strictly Increasing function.